• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:08:54 on December31,2007

    Entry number 222443


    keyword=Detector gas system alarm.

    There were detector gas system alarm since this morning:
    I checked the gas bottles, seems the problem was with ARGON
    gas: one bottle was full (1555 psi) while the other was around 400 psi. It did not switch automatically. I managed to get it switch and the alarms are gone. Argon now reads 1555 psi on the gui.
    However, the right VDC flow is still 0, need to to be checked.
    Alexandre Camsonne turned off the HV for right VDC earlier.

    All other gases are fine, including the N2 for NaI.

    On other thing I noticed: right NMR seems stuck. Will need to be rebooted.

    FIGURE 1